OpenStudy (anonymous):

Find the value of two numbers if their sum is 12 and their difference is 4

OpenStudy (anonymous):

guess and check for this one

OpenStudy (anonymous):

its system of operations

OpenStudy (anonymous):

x + y = 12 x - y = 4 2x = 16 x = 8 8 + y = 12 y = 4

OpenStudy (anonymous):

2 + 10 doesn't work because their difference is 8 3 + 9 doesn't work either because their difference is 6 try 4 and 8

OpenStudy (anonymous):

yeah x=8 and y =4

OpenStudy (mathmate):

When the sum and difference are known, the numbers are very easy to find. Larger number = (sum+difference)/2 = (12+4)/2 = 8 Smaller number = (sum-difference)/2 = (12-4)/2 = 4 So the two numbers are 8 and 4.